# LeoTCP: Low-Latency and High-Throughput Data Transport for LEO Satellite Networks

### v1

@@ III-C LeoTCP Sender
Reacting to path changes. To maintain stability when path changes and handovers occur, LeoTCP is designed to respond rapidly to RTT and congestion changes, without overreacting to packet loss events. Unlike traditional loss-based approaches [14, 39], which reduce the cwnd after three duplicate ACKs, LeoTCP adjusts its window solely based on INT telemetry, as detailed in Section III-C. This enables the sender to distinguish between congestive and non-congestive loss events. For retransmission and loss recovery, LeoTCP employs the TCP SACK extension [28] together with RACK [30]. SACK enables precise identification of lost packets, while RACK uses timers instead of packet ordering to detect losses, allowing LeoTCP to be more resilient to reordering. To mitigate the impact of reordering due to path changes, LeoTCP leverages the timestamp fields embedded in INT headers to ensure that only current per-hop telemetry is used for congestion estimation. Additionally, LeoTCP uses the pathID field (Section III-B) to detect path changes, allowing it to prioritise ACKs associated with the new path while safely ignoring outdated ACKs from old routes. To detect path changes, senders compare the XORed pathID from incoming ACKs to their currently stored value. If a new path is detected, the sender updates its stored pathID, but only once per sRTT interval to avoid reacting to reordering. If no change is detected, the existing value is maintained.
IV Experimental Evaluation
In this section, we evaluate the performance of LeoTCP using packet-level simulations conducted in OMNeT++ [19] and the INET framework [20]. We implemented Cubic, BBRv1, BBRv3, and our proposed LeoTCP protocol within OMNeT++/INET. We also updated INET’s TCP stack by refining the SACK implementation to better reflect the Linux kernel counterpart and added support for the RACK loss recovery mechanism [30]. RACK is used by all protocols in our evaluation and reflects common practice in modern TCP variants. All source code has been made publicly available to support reproducibility and future research. [^3] https://figshare.com/s/e503de35a07cd6326e06 We experimentally validated our models by comparing their performance against their Linux kernel counterparts within an emulation-based testbed. Although minor discrepancies arise due to differences in the underlying TCP stack implementations, the results exhibit agreement in performance trends and protocol behaviour, confirming the fidelity of the developed models. While [8, 17] are recent CC schemes designed for LEO satellite networks, their primary innovation lies in handling handovers by freezing the congestion window. Outside of these scenarios, they behave like their underlying CC algorithm. As our evaluation focuses broadly on LEO-specific challenges beyond just handovers, we omit the protocols from experiments and provide insights where handovers are explicitly tested. Unless otherwise stated, all experiments are conducted using 100Mbps links, and the results are averaged over five independent runs.
IV-A Experimenting with a LEO Satellite Network Model
We first evaluate how LeoTCP and existing congestion control algorithms behave in a realistic LEO satellite network setting, using a packet-level simulation model of LEO satellite networks [^4] Current LEO simulation models[21, 11] do not support hard handovers at the user terminal. As a result, we design a separate experiment in Section IV-D to specifically evaluate this scenario. [21] that supports both the BP and ISL topologies. We simulate Starlink’s first shell while incorporating 100 Starlink ground station locations. To reflect dynamic satellite connectivity and evolving RTT conditions, routing paths are updated every 100ms. Buffer sizes are set to match the BDP of the longer RTT path in each experiment. We conduct two sets of experiments: in the first, we evaluate both fairness and delay of two flows that share the same path throughout the simulation; in the second, we explore fairness, average goodput, and delay between two flows that follow independent paths with different RTTs but intermittently converge on a shared bottleneck due to routing changes.
@@ IV-A Experimenting with a LEO Satellite Network Model
Pair Cubic BBRv1 BBRv3 LeoTCP
@@
Figure 3(c) presents the average goodput ratio for two flows on different paths that intermittently share a bottleneck. The experiment includes two pairs of paths: Pair 1, with paths between New York to London and New York to St. John’s, Canada, which share a bottleneck for approximately 40% of the simulation; and Pair 2, with paths between San Diego to New York and Lawrence, Kansas to New York, with about 70% bottleneck overlap. Cubic performs well in both pairs, despite its tendency to favour lower RTT flows that react more quickly to loss [41]. This is because the results reflect average goodput across the full simulation that includes long periods where flows do not share a bottleneck; this reduces the impact of the short-term unfairness on the goodput ratio. BBRv1 shows substantial unfairness, particularly in Pair 2, as the bottlenecks are shared for a longer period of time. This stems from BBRv1’s bandwidth estimation and pacing, which favours flows with larger in-flight data. BBRv3 improves fairness over BBRv1 through its updated probe down phase, where the pacing gain increases from 0.75 to 0.91, causing the flow to reduce its rate less aggressively and limiting the ability of lower RTT flows to dominate the bandwidth. The normalised delay for all protocols mirrors characteristics from the previous experiment, with LeoTCP achieves both minimal delays and excellent fairness by ensuring that all flows compute a unified utilisation value through its avgRTT mechanism when competing for bandwidth.
IV-B Responsiveness
We now turn our attention to the experiment we used to motivate LeoTCP in the Introduction. We experiment with a simulated LEO satellite network path, accommodating a single 300-second flow. To emulate a highly dynamic LEO satellite network environment, we update the bottleneck bandwidth and base RTT values every 15 seconds. These values are uniformly selected from ranges of 50 to 100 Mbps for bandwidth and 5 to 100ms for RTT. We select RTT values that reflect the typical characteristics of current LEO networks [5, 11]. Varying bandwidth simulates different levels of congestion, for example, due to transient hotspots. We set the buffer capacity to $1\times$ the mean BDP of the path. For each CC approach, the experiment is repeated 50 times, and we calculate the cumulative distribution of the average goodput, as in [42].
@@ IV-B Responsiveness
As shown in Figure 1, both BBRv1 (solid green line) and BBRv3 (solid red line) consistently underutilise the available bandwidth due to their reliance on a minimum RTT filter, causing slow reaction to path changes. Cubic adapts quickly to bandwidth changes through its cubic window growth function (solid blue line), but this comes at the cost of significant latency inflation due to buffer overfilling. In contrast, LeoTCP adapts to dynamic conditions using hop-level congestion signals, reacting to bandwidth changes with accuracy and minimal delay inflation (solid orange line). Its explicit 95% utilisation target, prevents persistent queue build-up and ensures consistently low latencies.
In the presence of non-congestive loss, Cubic’s goodput degrades dramatically because Cubic interprets all loss as a congestion signal, unnecessarily reducing its sending rate (dashed blue line in Figure 1(a)). BBRv1, which entirely ignores loss, avoids overreacting and performs similarly to the no-loss case (dashed green line). BBRv3 continuously probes for bandwidth unless loss exceeds its 2% threshold, however, it still reacts aggressively even to mild loss, leading to poor goodput performance (dashed red line). In contrast, LeoTCP remains robust in lossy environments because it relies on in-network telemetry rather than loss as a congestion signal.
IV-C Fairness
@@
Figure 4 shows the goodput ratios of the multiple bottleneck setup. Cubic demonstrates behaviour close to proportional fairness, which becomes more pronounced as buffer sizes increase. Since Cubic reduces its window in response to packet loss, the flow crossing multiple bottlenecks encounters more frequent reductions, as it is more likely to encounter loss across several points in the path. BBRv3 and BBRv1 both exhibit fairness closer to proportional fairness. Although BBRv1 and BBRv3 use different probing mechanisms, in this homogeneous flow setup, flows traversing a single bottleneck end up probing roughly three times more frequently than the multi-bottlenecked flow, contributing to a fairness pattern closer to proportional fairness. LeoTCP, on the other hand, maintains strong fairness in the presence of multiple bottlenecks. By explicitly tracking the number of active flows and computing per-hop utilisation estimates, LeoTCP enables balanced bandwidth sharing across all flows, leading to bandwidth allocations that approximate max-min fairness.
@@
We next examine inter-RTT fairness, where flows competing for the same bottleneck experience different RTTs. We study how RTT affects fairness in such settings. In our experiment, one flow has a fixed RTT of 20ms, while the second flow’s RTT is varied. The queue size is set to $1\times$ the BDP of the higher RTT flow. The first flow is active for 2000 RTTs. The second flow starts 500 RTTs after the first flow. To allow CC schemes time to converge, we measure goodput over the final 500 RTTs of the experiment. Figure 5 presents the resulting goodput ratios. Cubic’s RTT bias resurfaces strongly here, as the flows with lower RTTs dominate. The BBR variants experience unfairness as RTTs diverge. This is because the flow experiencing a higher RTT will push more bytes in flight as a result of measuring a higher BDP, subsequently claiming higher buffer occupancy, leading to unfairness [40]. BBRv3 offers slightly improved fairness, as its probe down phase dampens the aggressiveness of rate reductions, but it still does not fully resolve its RTT unfairness. LeoTCP maintains consistent fairness, with goodput ratios close to 1 regardless of RTT disparity. This mirrors its earlier performance using the LEO simulation model, where its consistent utilisation calculation enabled fair bandwidth sharing.
@@
We now assess intra-RTT fairness, where two flows with the same base RTT compete over a shared bottleneck, otherwise repeating the methodology described in the inter-RTT experiment. Figure 6 shows the results. The base RTT of the flows is varied, as shown on the x-axis. Cubic demonstrates strong fairness across all RTT values, similar to its behaviour in Section IV-A. Both BBRv1 and BBRv3 perform slightly worse than Cubic. Upon closely inspecting individual runs, we observe that flows of both protocols take a long time to converge after the second flow is started, in comparison to LeoTCP and Cubic. LeoTCP consistently achieves the best fairness, maintaining near-optimal goodput ratios across all tested base RTT values.
IV-D Handovers
As discussed previously, path changes are frequent in LEO networks due to satellite movement. Soft and hard handovers can cause traffic shifts, leading to packet loss and transient congestion. To examine protocol behaviour under these conditions, we conduct two experiments: one simulates soft, lossless handovers, and the other introduces hard handovers with brief disconnect periods that induce packet loss.
@@ IV-D Handovers
We simulate a soft handover scenario between a GS and two satellites along a LEO satellite network path. Ground station handovers typically enable near-seamless transitions with minimal packet loss [8], although packet reordering may still occur. To evaluate LeoTCP’s responsiveness to such events, we simulate a setup where two flows are initially routed over a 20ms RTT path. At 100 seconds, these flows are redirected to a different network path that already serves two flows with a higher RTT (50ms). At 200 seconds, the rerouted flows return to their original lower RTT path. To assess the impact of redirection, we measure the average normalised goodput and Jain’s fairness index after the path change (“Change 1”), and again after the flows return to their original paths (“Change 2”). Measurements are over the first 100 RTTs after each change to evaluate protocols’ ability to adapt quickly and fairly.
Cubic and the BBR variants show higher unfairness during the first change (triangles in Figure 7), consistent with the inter-RTT unfairness observed above. Following the second change at 200 seconds (Change 2), the experiment transitions into an intra-RTT setup, with both flow groups operating under their respective base RTTs, showing once again similar fairness to the intra-RTT experiment. LeoTCP maintains high fairness during both transitions. The high fairness during Change 1 is attributed to its avgRTT and INT-driven responsiveness. In terms of bandwidth exploration, Cubic does not capture much of the available bandwidth in Change 2, (blue star in Figure 7). This is attributed to its concave window growth period, which must elapse before Cubic can explore beyond its previously seen maximum congestion window. A higher queue size would allow Cubic to carry a much larger congestion window into Change 2, resulting in a faster bandwidth exploration, at the cost of a higher RTT. The BBR variants must enter the probe bandwidth phase to begin bandwidth exploration. BBRv1 is much quicker compared to BBRv3, as the former is RTT based and the latter is wall-clock-based. LeoTCP also maintains high goodput during both changes due to the INT information allowing senders to immediately determine the characteristics of the new path once the first ACK is received after the change. This is further enabled by its use of the pathID field, which allows the sender to accurately detect the path change and immediately prioritise feedback from the new route.
@@
In terms of latencies, Cubic’s RTT reaches 61.96ms for the rerouted flow on the 20ms base RTT path and 93.77ms for the constant flow on the 50ms base RTT path, indicating substantial queue buildup at the bottleneck. BBRv1 shows similar inflation, with RTTs of 63.94ms (20ms path) and 94.63ms (50ms path), consistent with earlier results in the LEO simulation model (Section IV-A), where BBRv1 tends to overestimate available bandwidth under multi-flow conditions. BBRv3 performs slightly better, recording 46.36ms (20ms path) and 77.77ms (50ms path) RTTs, due to it reacting to loss, as well as probing for bandwidth less frequently than BBRv1. LeoTCP exhibits the lowest delay inflation of all protocols, with RTTs of just 21.50ms (20ms path) and 51.50ms (50ms path). These closely track their respective base RTTs, as LeoTCP explicitly targets 95% utilisation, keeping queueing minimal even during path changes.
Hard Handovers
We next explore protocol performance under hard handovers, which result in abrupt interruptions and packet loss. To simulate this, we configure a single flow over a 50ms LEO satellite network path, where hard handovers are triggered every 5 seconds. Each handover introduces a complete path disconnection lasting between 20ms and 200ms (as shown on the x-axis). During these interruptions, all in-flight packets and any queued packets are discarded. We measure goodput for 50 RTTs following each disconnection event, allowing us to assess the performance degradation caused by varying interruption durations.
Figure 8: Goodput of a single flow with varying path interruptions (shown on the x-axis) every 5 seconds.
Cubic suffers greatly under these conditions, as each handover-induced loss triggers a multiplicative decrease of its congestion window. When the disruption period increases, all protocols experience retransmission timeouts. For Cubic, this timeout triggers a slow start phase, allowing Cubic to rapidly regain bandwidth. BBRv1 and BBRv3 also experience performance degradation when their probing coincides with a disruption. Since BBR relies on periodic probing to update its model, a failed probe during a handover can cause it to reduce its sending rate and stall until the next scheduled probe, further limiting goodput. LeoTCP recovers from each disruption swiftly. By relying on INT and not using packet loss as a congestion signal, it maintains accurate congestion estimates and resumes sending at the previously established rate once the path is restored. Although not evaluated in this experiment, [8, 17] are designed to freeze their states during disruption periods, avoiding reductions from handover loss, therefore mitigating their underlying CCs throughput degradation. We would therefore expect those approaches to perform similarly to LeoTCP.
IV-E LeoTCP Design Evaluation
This section evaluates two key design components of LeoTCP that enable high performance in LEO satellite networks: (1) the use of a per-switch avgRTT to ensure fair utilisation across flows with varying propagation delays, and (2) the initial phase mechanism, which enables faster convergence during flow start-up, particularly when many flows contend for the same bottleneck.
Average RTT
@@ IV-E LeoTCP Design Evaluation
Initial Phase
@@
We now evaluate the effectiveness of LeoTCP’s initial phase mechanism in enabling fast convergence for new flows in a shared bottleneck scenario. The experiment uses a dumbbell topology, where all flows traverse the same bottleneck link with a fixed RTT of 100ms and a bandwidth of 200Mbps. Initially, 50 flows start within a 5 RTT window. Additional flows then join incrementally: 1 new flow at 10s, 2 more at 20s, 4 at 30s, and 8 at 40s. Figure 10 shows the evolution of goodput over time for all flows. This setup tests whether LeoTCP’s initial phase design allows newly arriving flows to quickly ramp up to their fair share without affecting the goodput of existing flows.
@@
As shown in Figure 10, without the initial phase, new flows take a large amount of RTTs to reach their fair share of bandwidth. With the initial phase enabled, LeoTCP waits one avgRTT after receiving INT feedback. This allows the sender to obtain an accurate estimate of the number of competing flows before increasing its rate. As a result, newly arriving flows ramp up smoothly and quickly, without disrupting existing traffic. Despite being more aggressive during start-up, utilising the initial phase introduces only minimal additional queuing. The average RTT of the first 50 flows, measured over 100 RTTs at the start of the simulation, is 113.50ms with the initial phase enabled, compared to 103.05ms without. When eight additional flows join at 40 seconds, the difference is less, with a 105.53ms RTT with the initial phase, versus 104.18ms without. These results confirm that LeoTCP can achieve rapid convergence at startup without significantly compromising latency, as both new and existing flows can quickly adapt to changes in network conditions using INT.
V Conclusion

### v2

@@ II OrbCC Design
Reacting to path changes. To maintain stability when path changes and handovers occur, OrbCC is designed to respond rapidly to RTT and congestion changes, without overreacting to packet loss events. Unlike loss-based approaches [15, 11], which reduce the cwnd after three duplicate ACKs, OrbCC adjusts its window solely based on the provided in-network information. This enables the sender to distinguish between congestive and non-congestive loss events. For retransmission and loss recovery, OrbCC employs the TCP SACK extension [10] together with RACK [7]. SACK enables precise identification of lost packets, while RACK uses timers instead of packet ordering to detect losses, allowing OrbCC to be more resilient to reordering. To mitigate the impact of reordering due to path changes, OrbCC leverages the timestamp fields embedded in its headers to ensure that only current network provided information is used for congestion estimation. Additionally, OrbCC uses the pathID field to detect path changes, allowing it to prioritise ACKs associated with the new path while safely ignoring outdated ACKs from old routes. To detect path changes, senders compare the XORed pathID from incoming ACKs to their currently stored value. If a new path is detected, the sender updates its stored pathID, but only once per sRTT interval to avoid reacting to reordering. If no change is detected, the existing value is maintained.
III Experimental Evaluation
We evaluate OrbCC using packet-level simulations in OMNeT++ [32] and INET [19], extended with a LEO satellite model [^3] Current LEO simulation models[35, 22] do not support hard handovers at the user terminal. Therefore, we design separate experiments in Sections III and $h$ to specifically evaluate these scenarios. [35] supporting BP and ISL topologies. Our setup simulates Starlink’s first shell with 100 ground stations, updating routing every 100ms to reflect dynamic connectivity and RTT variation. Buffer sizes match the BDP of the longer RTT path. We implemented Cubic, BBRv1, BBRv3, SaTCP, LeoCC and OrbCC refining INET’s SACK implementation to better reflect the Linux kernel counterpart and added support for the RACK loss recovery mechanism [7]. RACK reflects common practice in modern TCP variants. Source code is available. [^4] https://figshare.com/s/334dbee6891e4edf6f9e We validated our models against their Linux kernel counterparts within an emulation-based testbed, showing agreement in performance trends and protocol behaviour. SaTCP and LeoCC both modify CC behaviour in response to handovers; outside of these events it behaves like its underlying CC. We therefore include SaTCP and LeoCC only in experiments that explicitly model hard handovers. We note that the original SaTCP design assumes user-terminals provide live information on its handovers. In our simulations, handover events are known and provided to the protocol. Unless stated, links are 100Mbps and results average five runs.
Using the LEO constellation model [35], we first examine fairness and delay when two flows share the same path, and then when flows traverse different RTT paths that intermittently converge on a bottleneck. Since the model does not capture hard handovers, we evaluate Cubic, BBRv1, BBRv3, and OrbCC in these experiments.
@@ III Experimental Evaluation
Pair Cubic BBRv1 BBRv3 OrbCC
@@
Figure 3(c) presents the average goodput ratio for two flows on different paths that intermittently share a bottleneck. The experiment includes two pairs of paths: Pair 1, with paths between New York to London and New York to St. John’s, Canada, which share a bottleneck for approximately 40% of the simulation; and Pair 2, with paths between San Diego to New York and Lawrence, Kansas to New York, with about 70% bottleneck overlap. Cubic performs well in both pairs, despite its tendency to favour lower RTT flows that react more quickly to loss [12]. This is because the results reflect average goodput across the full simulation that includes long periods where flows do not share a bottleneck; this reduces the impact of the short-term unfairness on the goodput ratio. BBRv1 shows substantial unfairness, particularly in Pair 2, as the bottlenecks are shared for a longer period of time. This stems from BBRv1’s bandwidth estimation and pacing, which favours flows with larger in-flight data. BBRv3 slightly improves fairness over BBRv1 through its updated probe down phase, where the pacing gain increases from 0.75 to 0.91, causing the flow to reduce its rate less aggressively and limiting the ability of lower RTT flows to dominate the bandwidth. The normalised delay for all protocols mirrors characteristics from the previous experiment, with OrbCC achieving both lower delays and excellent fairness by ensuring that all flows compute a unified utilisation value through its avgRTT mechanism when competing for bandwidth.
Responsiveness
We revisit the motivating experiment from the Introduction: a single flow over a LEO path with bandwidth and RTT changing every 5s, where each reconfiguration induces a hard handover outage of 45-120ms. Bandwidth varies from 50–100Mbps and RTT from 1–100ms, with buffers set to the mean BDP. Each protocol is run 50 times.
@@
As shown in Figure 1, both BBRv1 (solid green line) and BBRv3 (solid red line) consistently underutilise the available bandwidth due to their reliance on a minimum RTT filter, causing slow reaction to path changes. Cubic adapts quickly to bandwidth changes through its cubic window growth function (solid blue line), but this comes at the cost of significant latency inflation due to buffer overfilling. Handover-induced losses repeatedly trigger Cubic’s multiplicative backoff, leading to sustained underutilisation. SaTCP, by freezing its congestion window across reconfigurations, avoids this handover-driven underutilisation, but still incurs large delay due to buffer build-up (solid purple line). LeoCC attains similarly high goodput due to it resetting its model and entering startup post-reconfiguration (solid teal line). However, LeoCC’s startup exponentially ramps-up until a stable bandwidth is observed for three consecutive RTTs, inflating queues and inducing loss. In contrast, OrbCC reacts to reconfigurations using its per-hop congestion information, tracking bandwidth changes accurately with minimal delay inflation (solid orange line). Its explicit 95% utilisation target prevents persistent queue build-up and ensures consistently low latencies.
In the presence of non-congestive loss, both Cubic and SaTCP’s goodput degrades dramatically because both CCs interpret loss as a congestion signal, unnecessarily reducing its sending rate (dashed blue and purple line in Figure 1(a)). Both BBRv1 and its derivative, LeoCC, ignore loss and therefore avoid overreacting, performing similarly to the no-loss cases (dashed green and teal lines). BBRv3 continuously probes for bandwidth unless loss exceeds its 2% threshold, however, it still reacts aggressively even to mild loss, leading to poor goodput performance (dashed red line). OrbCC remains robust in lossy environments as it does not use loss as a congestion signal.
@@
Figure 4 shows the goodput ratios of the multiple bottleneck setup. Cubic demonstrates behaviour close to proportional fairness, which becomes more pronounced as buffer sizes increase. Since Cubic reduces its window in response to packet loss, the flow crossing multiple bottlenecks encounters more frequent reductions, as it is more likely to encounter loss across several points in the path. BBRv3 and BBRv1 both exhibit fairness closer to proportional fairness. Although BBRv1 and BBRv3 use different probing mechanisms, in this homogeneous flow setup, flows traversing a single bottleneck end up probing roughly three times more frequently than the multi-bottlenecked flow, contributing to a fairness pattern closer to proportional fairness. OrbCC, on the other hand, maintains strong fairness in the presence of multiple bottlenecks. By explicitly tracking the number of active flows and computing per-hop utilisation estimates, OrbCC enables balanced bandwidth sharing across all flows, leading to bandwidth allocations that approximate max-min fairness.
@@
We next examine inter-RTT fairness, where flows competing for the same bottleneck experience different RTTs. One flow has a fixed RTT of 20ms, while the second flow’s RTT is varied. The queue size is set to $1\times$ the BDP of the higher RTT flow. The first flow is active for 2000 RTTs. The second flow starts 500 RTTs after the first flow. To allow CC schemes time to converge, we measure goodput over the final 500 RTTs of the experiment. Figure 5 presents the resulting goodput ratios. Cubic’s RTT bias resurfaces strongly here, as the flows with lower RTTs dominate. The BBR variants experience unfairness as RTTs diverge. This is because the flow experiencing a higher RTT will push more bytes in flight as a result of measuring a higher BDP, subsequently claiming higher buffer occupancy, leading to unfairness [34]. BBRv3 offers slightly improved fairness, as its probe down phase dampens the aggressiveness of rate reductions, but it still does not fully resolve its RTT unfairness. OrbCC maintains consistent fairness, with goodput ratios close to 1 regardless of RTT disparity. This mirrors its earlier performance using the LEO simulation model, where its consistent utilisation calculation enabled fair bandwidth sharing.
@@
We now assess intra-RTT fairness, where two flows with the same base RTT compete over a shared bottleneck, otherwise repeating the methodology described in the inter-RTT experiment. In Figure 6, the base RTT of the flows is varied, as shown on the x-axis. Cubic demonstrates strong fairness across all RTT values, similar to its behaviour in Section III. Both BBRv1 and BBRv3 perform slightly worse than Cubic. Upon closely inspecting individual runs, we observe that flows of both protocols take a long time to converge after the second flow is started, in comparison to OrbCC and Cubic. OrbCC consistently achieves the best fairness, maintaining near-optimal goodput ratios across all tested base RTT values.
Soft and hard handovers can cause traffic shifts, leading to packet loss and transient congestion. We conduct one simulation of soft, lossless handovers, and the other looks further into hard handovers following Section III, evaluating how increasing the number of concurrent flows affects aggregate goodput and delay under hard handover disruptions.
@@
We simulate a soft handover scenario between a GS and two satellites along a LEO satellite network path. Ground station handovers typically enable near-seamless transitions with minimal packet loss [5], although packet reordering may still occur. To evaluate OrbCC’s responsiveness to such events, two flows are initially routed over a 20ms RTT path. At 100 seconds, these flows are redirected to a different network path that already serves two flows with a higher RTT (50ms). At 200 seconds, the rerouted flows return to their original lower RTT path. To assess the impact of redirection, we measure the average normalised goodput and Jain’s fairness index after the path change (“Change 1”), and again after the flows return to their original paths (“Change 2”). Measurements are taken over the first 100 RTTs after each change to evaluate the protocols’ ability to adapt quickly and fairly.
@@
Cubic and the BBR variants show higher unfairness during the first change (triangles in Figure 7), consistent with the inter-RTT unfairness observed above. Following the second change at 200 seconds (Change 2), the experiment transitions into an intra-RTT setup, with both flow groups operating under their respective base RTTs, showing once again similar fairness to the intra-RTT experiment. OrbCC maintains high fairness during both transitions. The high fairness during Change 1 is attributed to its avgRTT and switch-reported path signals. In terms of bandwidth exploration, Cubic does not capture much of the available bandwidth in Change 2, (blue star in Figure 7). This is attributed to its concave window growth period, which must elapse before Cubic can explore beyond its previously seen maximum congestion window. A higher queue size would allow Cubic to carry a much larger congestion window into Change 2, resulting in a faster bandwidth exploration, at the cost of a higher RTT. The BBR variants must enter the probe bandwidth phase to begin bandwidth exploration. BBRv1 is much quicker compared to BBRv3, as the former is RTT based and the latter is wall-clock-based. OrbCC also maintains high goodput during both changes due to it allowing senders to immediately determine the characteristics of the new path once the first ACK is received after the change. This is further enabled by its use of the pathID field, which allows the sender to accurately detect the path change and immediately prioritise feedback from the new route.
Averaged over 100 RTTs following the first path change, Cubic’s RTT reaches 62.7ms for the rerouted flow on the 20ms base RTT path and 94.57ms for the constant flow on the 50ms base RTT path, indicating substantial queue buildup at the bottleneck. BBRv1 shows similar inflation, with RTTs of 63.45ms (20ms path) and 92.44ms (50ms path), consistent with earlier results in the LEO simulation model (Section III), where BBRv1 tends to overestimate available bandwidth under multi-flow conditions. BBRv3 performs slightly better, recording 43.63ms (20ms path) and 78.74ms (50ms path) RTTs, due to it reacting to loss, as well as probing for bandwidth less frequently than BBRv1. OrbCC exhibits the lowest delay inflation of all protocols, with RTTs of just 21.50ms (20ms path) and 51.50ms (50ms path). These closely track their respective base RTTs, as OrbCC explicitly targets 95% utilisation, keeping queueing minimal even during path changes.
Hard Handover Efficiency
@@
Figure 8: Normalised goodput vs normalised delay of multiple flows sharing the bottleneck
We now evaluate two key design components of OrbCC that enable high performance in LEO satellite networks: (1) the use of a per-switch avgRTT to ensure fair utilisation across flows with varying propagation delays, and (2) the initial phase mechanism, which enables faster convergence during flow start-up, particularly when many flows contend for the same bottleneck.
@@
OrbCC Average RTT
[…]
OrbCC Initial Phase
@@
We now evaluate the effectiveness of OrbCC’s initial phase mechanism in enabling fast convergence for new flows in a shared bottleneck scenario. The experiment uses a dumbbell topology, where all flows traverse the same bottleneck link with a fixed RTT of 100ms and a bandwidth of 200Mbps. Initially, 50 flows start within a 5 RTT window. Additional flows then join incrementally: 1 new flow at 10s, 2 more at 20s, 4 at 30s, and 8 at 40s. Figure 10 shows the evolution of goodput over time for all flows. This setup tests whether OrbCC’s initial phase design allows newly arriving flows to quickly ramp up to their fair share without affecting the goodput of existing flows.
@@
As shown in Figure 10, without the initial phase, new flows take a large amount of RTTs to reach their fair share of bandwidth. With the initial phase enabled, OrbCC waits one avgRTT after receiving in-network information. This allows the sender to obtain an accurate estimate of the number of competing flows before increasing its rate. As a result, newly arriving flows ramp up smoothly and quickly, without disrupting existing traffic. Despite being more aggressive during start-up, utilising the initial phase introduces only minimal additional queuing. The average RTT of the first 50 flows, measured over 100 RTTs at the start of the simulation, is 104.68ms with the initial phase enabled, compared to 104.13ms without. When eight additional flows join at 40 seconds OrbCC has an average RTT of 105.48ms RTT with the initial phase, versus 104.18ms without. These results confirm that OrbCC can achieve rapid convergence at startup without significantly compromising latency, as both new and existing flows can quickly adapt to changes in network conditions.
IV Discussion and Conclusion
OrbCC targets operator-managed LEO constellations under a single administrative domain [29], enabling uniform deployment across satellites and user terminals; we therefore designed fairness among competing OrbCC flows. This deployment model is consistent with other LEO approaches [5, 25] that assume gateway/middlebox assistance, making TCP-friendliness within the satellite segment a secondary concern. In this setting, performance-enhancing proxies (PEPs) can also be practical [39]: gateways can split or proxy connections at the constellation edge, isolating terrestrial endpoints from LEO-specific dynamics and allowing OrbCC to operate optimally within the satellite segment while preserving end-to-end semantics where required. Compatibility with conventional end-to-end CC is primarily needed at terrestrial interconnects, where gateways can hand off or translate between OrbCC within the constellation and an end-host CC approach outside it.
The current design of OrbCC assumes that its in-network signalling mechanisms are available at all intermediate routers along the path. Under a partial or incremental deployment of OrbCC switches, packets would carry incomplete network state information, potentially limiting the sender’s ability to accurately locate bottlenecks and distinguish between propagation and queueing delays. We leave the design and evaluation of mechanisms for robust operation in such environments as future work, but foresee that OrbCC could fall back to utilising alternate state-of-the-art LEO CC approaches [25].